### v1

# Machine Learning in Fundamental Physics Natural Image Classification via Quasi-Cyclic Graph Ensembles and Random-Bond Ising Models at the Nishimori Temperature

Abstract
We propose a unified framework that brings together statistical physics, coding theory, and algebraic topology for multi-class image classification. First, each input image is passed through a frozen MobileNetV2 backbone to obtain a high-dimensional feature vector $\mathbf{h}\in\mathbb{R}^{1280}$. The components of $\mathbf{h}$ are interpreted as Ising spins placed on the vertices of a sparse Multi-Edge Type quasi-cyclic LDPC (MET-QC-LDPC) graph. The resulting Random-Bond Ising Model (RBIM) is operated at its Nishimori temperature $\beta_{N}$, which we identify as the unique point where the smallest eigenvalue of the Bethe–Hessian matrix $H_{\beta,J}$ vanishes: $\lambda_{\min}\!\bigl{(}H_{\beta_{N},J}\bigr{)}=0.$ At this critical temperature class separability is maximized. Our first theoretical contribution is an exact correspondence between local trapping sets in the Tanner graph of a QC-LDPC code and topological invariants of the underlying feature manifold. We prove that each elementary cycle generates a pole of the Ihara–Bass zeta function, which appears as an isolated eigenvalue of both the non-backtracking operator and $H_{\beta,J}$. These spectral signatures encode $\mathbb{Z}/2$-torsion (Betti numbers, bordism classes) that act as topological defects in the embedding. The second contribution is a practical algorithm for estimating $\beta_{N}$. By evaluating $\lambda_{\min}(H_{\beta,J})$ at three trial temperatures, fitting a quadratic interpolant, and applying one Newton correction, the method converges in roughly nine Arnoldi iterations—approximately a six-fold speed-up over standard bisection. Guided by the topological analysis we design two MET-QC-LDPC graph ensembles (a spherical family with a single circulant ring and a toroidal family with two independent rings). Permanent and Bethe–permanent bounds are used to suppress harmful trapping sets during protograph and parity-check matrix optimization. The resulting spectral embeddings compress the original 1280-dimensional MobileNetV2 features to $32$ dimensions for a 10-class ImageNet subset (ImageNet-10), and $64$ dimensions for a 100-class ImageNet subset (ImageNet-100). Despite this compression we achieve 98.7% top-1 accuracy on ImageNet-10 (spherical graphs) and 82.7% top-1 accuracy on ImageNet-100 using a three-graph ensemble, while reducing the number of parameters and memory footprint by factors of $40$ and $20$, respectively. In summary, controlling graph design through topological invariants yields highly efficient, physics-inspired embeddings that retain state-of-the-art classification performance.
Clustering, Classification, Cobordism, Feature Embeddings, Graph codes, Multi-Edge Type graphs, Nishimori, Random Bond Ising Model, Quasi-Cyclic, QC-LDPC, Spectral Gap, Stochastic Block Model
pacs:
02.10.Ox, 02.10.Yn, 02.40.Vh, 02.40.Re, 02.40.-k, 02.70.Hm, 03.65.Db, 05.70.Ln, 64.70.Tg
1 Introduction
Graph-based spectral methods have recently shown great promise for machine learning tasks. In particular, modeling feature vectors as spins (nodes) on a sparse graph under a Gibbs distribution (a Random-Bond Ising Model) enables powerful clustering and classification. For example, Nishimori and Bethe free-energy ideas have been used to tune graph-based classifiers ([1] , and prior work demonstrated that sparse QC-LDPC graphs with Nishimori-weighted edges can significantly improve clustering of high-dimensional CNN features, [2]. However, most of these successes have been in synthetic or binary settings (e.g. distinguishing GAN-generated images), while natural multi-class image datasets pose additional challenges: the feature manifolds can have nontrivial topology and curvature that corrupt spectral embeddings, and standard methods struggle to capture nonlinear class boundaries. In this work, we develop a rigorous, end-to-end framework that unifies ideas from statistical physics, coding theory, and topology to tackle these challenges. Our contributions are threefold. We derive explicit relationships between graph-theoretic trapping sets (small subgraphs in Tanner graphs) and spectral/topological invariants. In particular, we link closed cycles in the graph to zeros and poles of the Ihara–Bass zeta function, and show how these induce isolated eigenvalues in the Bethe–Hessian and non-backtracking operators. This analysis reveals that trapping sets act as torsion defects in the feature manifold, which can be characterized by topological invariants (Betti numbers, quadratic-form signature, etc.) and by $K$-theory indices. We introduce a novel algorithm to estimate the Nishimori temperature in graph-based RBIMs. By fitting a low-degree Lagrange interpolant to the smallest Bethe–Hessian eigenvalue and imposing a stability constraint, our method converges reliably to the critical $\beta_{N}$ even under complex geometric constraints. This yields a provably optimal spin-glass/paramagnetic threshold for clustering. We apply ensembles of three expert quasi-cyclic LDPC graphs to real image classification benchmarks (ImageNet-10 and ImageNet-100). Starting from 1280-dim MobileNetV2 features, we use these structured sparse graphs to perform a spectral embedding that compresses features to 32 (resp. 64) dimensions per class. Using our Nishimori-tuned RBIM weights and ensemble majority voting, we achieve 98.7% accuracy on ImageNet-10 and 82.7% on ImageNet-100 (comparable to a 64-D dense embedding), while reducing parameters by a factor of 20×. The remainder of the paper is organised as follows. Section 2 recalls the RBIM and the Nishimori condition. Section 3 introduces quasi-cyclic LDPC graphs and their multi-edge extensions. Topological consequences of trapping sets are discussed in Section 4. The temperature estimator is presented in Section 4.6. Experimental results on ImageNet are reported in Section 5, and we conclude with future directions in Sections  6,  Conclusion.
2 Random Bond Ising Models
Let $\mathcal{G}=(V,E)$ be an undirected graph with $|V|=n$. Assign a binary spin variable $s_{i}\in\{-1,+1\}$ to each vertex $i\in V$ and denote the spin configuration by $\mathbf{s}=(s_{1},\dots,s_{n})^{\!\top}$. For a symmetric coupling matrix $J=(J_{ij})_{i,j=1}^{n}$, the Hamiltonian of the Random-Bond Ising Model is
$\mathcal{H}_{J}(\mathbf{s})=-\sum_{i<j}J_{ij}s_{i}s_{j}=-\tfrac{1}{2}\,\mathbf{s}^{\!\top}J\mathbf{s}.$
At inverse temperature $\beta$ the Boltzmann distribution:
$\mu_{\beta,J}(\mathbf{s})=\frac{\exp\bigl{(}-\beta\,\mathcal{H}_{J}(\mathbf{s})\bigr{)}}{Z_{J,\beta}},\qquad Z_{J,\beta}=\sum_{\mathbf{s}\in\{-1,+1\}^{n}}\!e^{-\beta\mathcal{H}_{J}(\mathbf{s})}.$
In a classification setting the coupling $J$ encodes similarity between feature vectors:
$J_{ij}=\begin{cases}+1,&\text{if }y_{i}=y_{j},\\ -1,&\text{otherwise},\end{cases}$
where $y_{i}$ denotes the class label of vertex $i$. More refined couplings are obtained from a weighted graph built on the CNN features (see Section 4.7).
2.1 Nishimori condition and phase transition
For a disordered Ising system the Nishimori line [3, 4] specifies a particular temperature $\beta_{N}$ at which the distribution of couplings matches that of the spins. If the edge weights $J_{ij}$ are drawn from a symmetric distribution $P(J)$ satisfying
$P(J)=p(|J|)\,e^{\beta_{N}J},\qquad\int p(|J|)e^{\beta_{N}J}\,\mathrm{d}J=1,$
the Nishimori condition holds. At $\beta=\beta_{N}$ the model exhibits a paramagnetic–spin-glass transition, and many observables (e.g. magnetisation, overlap) become analytically tractable.
A convenient characterisation of $\beta_{N}$ uses the Bethe–Hessian matrix $H_{\beta,J}$ (see Section 2.2). Its smallest eigenvalue $\lambda_{\min}(\beta)$ is a strictly decreasing function of $\beta$, and the Nishimori point is uniquely defined by
$\lambda_{\min}(\beta_{N})=0.$ (1)
Thus estimating $\beta_{N}$ reduces to solving a one-dimensional root–finding problem.
2.2 Bethe free energy and the Bethe–Hessian
Let $q_{i}=\Pr(s_{i}=+1)$ be an approximate marginal. The Bethe free energy associated with a factorized distribution $p_{q}(\mathbf{s})=\prod_{i}q_{i}^{(1+s_{i})/2}(1-q_{i})^{(1-s_{i})/2}$,  [6]:
$\widetilde{F}_{J,\beta}(q)=\sum_{\mathbf{s}}p_{q}(\mathbf{s})\bigl{[}\,\beta\,\mathcal{H}_{J}(\mathbf{s})+\ln p_{q}(\mathbf{s})\bigr{]}.$
The Hessian of $\widetilde{F}_{J,\beta}$ with respect to the vector $\mathbf{m}=(2q_{i}-1)_{i}$ is the Bethe–Hessian (also called deformed Laplacian)  [7, 1]:
$H_{\beta,J}=\operatorname{diag}\!\Bigl{(}1+\sum_{k\in\partial i}\frac{\tanh^{2}(\beta J_{ik})}{1-\tanh^{2}(\beta J_{ik})}\Bigr{)}-\biggl{[}\frac{\tanh(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}\biggr{]}_{i\neq j}.$ (2)
At the Nishimori temperature $\beta_{N}$ the smallest eigenvalue of $H_{\beta,J}$ vanishes, signalling a critical point where community structure becomes most pronounced, Fig. 1.
Figure 1: In (a) and (b) we see cluster structures (community) from phase transaction described by adjacency matrices and network related graph, Fig. 20 from [5]
3 Multi-Edge Type Quasi-Cyclic LDPC Graphs
Low-density parity-check (LDPC) codes naturally define sparse bipartite graphs. A Quasi-Cyclic LDPC (QC-LDPC) codes provide a structured and hardware-friendly subclass defined by a quasi-cyclic parity-check matrix $H$ [8]. An $(N,K)$ QC-LDPC code consists of $N$ total codeword bits, with $K$ information bits and NKN-K parity bits. The associated Tanner graph is described by a parity-check matrix $H\in\mathbb{F}_{2}^{mL\times nL}$, constructed from square blocks of size $L\times L$, where each block is either a zero matrix or a circulant permutation matrix (CPM) [9, 10].
A CPM $P\in\{0,1\}^{L\times L}$ is defined as:
$P_{ij}=\begin{cases}1,&\text{if }i+1\equiv j\pmod{L},\\ 0,&\text{otherwise}.\end{cases}$
Let $P_{k}$ denote a circulant permutation matrix corresponding to a right shift of the identity matrix $I$ by $k\in\{0,1,\dots,L-1\}$, ring of size $\mathbb{Z}/L\mathbb{Z}$. Then, a general parity-check matrix $H_{\text{QC}}$ of a QC-LDPC code takes the block form:
$H_{\text{QC}}=\begin{bmatrix}P_{a_{11}}&P_{a_{12}}&\dots&P_{a_{1n}}\\ P_{a_{21}}&P_{a_{22}}&\dots&P_{a_{2n}}\\ \vdots&\vdots&\ddots&\vdots\\ P_{a_{m1}}&P_{a_{m2}}&\dots&P_{a_{mn}}\end{bmatrix}$,
where each $a_{ij}\in\mathcal{A}_{L}=\{0,1,\dots,L-1\}$. The circulant size $L$ controls both the code length and the degree of quasi-cyclicity. When the parity-check matrix consists of quasi-cyclic rings of circulants, we refer to this configuration as a toroidal graph family. As an example of toroidal graph, consider a QC-LDPC code defined by the parity-check matrix, [11]:
$H=\begin{bmatrix}I_{1}&I_{2}&I_{4}\\ I_{6}&I_{5}&I_{3}\\ \end{bmatrix}$,
of size $31\times 21$ with circulant size $L=7$, employing a two-ring (3 columns and 2 rows) construction $(\mathbb{Z}/7\mathbb{Z})$ as illustrated in Fig. 1 (left).
From $H_{\text{QC}}$, two key matrices can be derived, exponent matrix $E(H)$, containing shift values $a_{ij}$, the protograph matrix $M(H_{\text{QC}})$, where each nonzero CPM is replaced by 1 and zeros are left as is.
Figure 2: (Left) Tanner graph corresponding to the parity-check matrix $H_{1}$, composed by 2 and 3 ring of size 7, Fig.2 [11]. (Right) Multi-graph protograph $M(H_{2})$ representation of the QC-LDPC code $H_{2}$.
An illustrative example of a protograph representation is given in Fig. 2. Consider the following QC-LDPC matrix:
$H_{2}=\begin{bmatrix}I_{1}+I_{2}+I_{7}&I_{9}&I_{23}&0&0\\ I_{12}+I_{37}&I_{19}&0&I_{32}&I_{11}+I_{12}\\ 0&0&I_{33}&0&0\end{bmatrix}$,
where $I_{k}$ denotes a CPM corresponding to a shift of $k$. The use of CPM sums in $H_{2}$ characterizes a Multi-Edge Type (MET) QC-LDPC code [12, 13], allowing for more flexible degree distributions and improved code performance under iterative decoding. Family of (MET) QC-LDPC graph codes due two (independent) ring of circulant we shall call toroidal graphs.
In the special case where the parity check matrix contains only one circulant ring with multiple weights, we shall call it a spherical graph family. For example, such a parity-check matrix with CPM size $L=2600$ is represented below, [14, 15]:
$H_{3}=\begin{bmatrix}I_{2}+I_{3}+I_{5}+I_{280}+I_{437}+I_{511}+I_{636}+I_{797}\\ +I_{1022}+I_{1093}+I_{1233}+I_{1671}+I_{1718}+I_{2254}+I_{2334}\end{bmatrix}$.
[…]
$A=\begin{bmatrix}0&H_{\text{QC}}^{\!\top}\\ H_{\text{QC}}&0\end{bmatrix},\qquad D=\operatorname{diag}(A\mathbf{1}).$
This adjacency will serve as the basis for the coupling matrix $J$ in the RBIM.
4 Topological Signatures of Trapping Sets
Trapping sets (TS) in Tanner graphs correspond to local topological defects that disrupt the decoding dynamics of LDPC codes, [16]. A trapping set $TS(a,b)$, formed by cycles (block-cycle for QC-LDPC) or cycle (block-cycle) overlap, consists of $a$ variable nodes and $b$ odd-degree check nodes, where the configuration prevents successful iterative decoding. When $b=0$, such sets correspond to codewords, i.e., $TS(a,0)$ with $a=d_{\min}$, the minimum codeword weight, [16]. To evaluate the harmfulness of cycles, the Extrinsic Message Degree (EMD) metric is used, [17]. It quantifies the number of singly connected check nodes to a given cycle. For practical computation, the Approximate Cycle EMD (ACE) is employed [17]: $\mathrm{ACE}(C)=\sum_{v\in E(V_{c})}\left(d(v)-2\right)$, where $C$ is a cycle in the graph, $d(v)$ is the degree of a variable node $v$, and $V_{c}$ denotes the set of variable nodes within cycle $C$. By enforcing higher minimum ACE values for cycles of fixed length, one mitigates harmful TS. Example of TS(4,2), TS(4,6), TS(9,2) defined by matrices (3), represented on Fig. 3, $x_{i}$ - columns, $c_{i}$ - rows of $H_{TS}$.
$H_{TS(4,2)}\begin{bmatrix}1&0&0&0\\ 1&1&0&0\\ 0&1&1&0\\ 0&0&1&1\\ 0&0&0&1\end{bmatrix},H_{TS(4,6)}=\begin{bmatrix}0&1&0&0\\ 1&0&0&0\\ 1&1&0&0\\ 0&1&0&0\\ 1&0&0&0\\ 1&1&0&0\\ 0&1&0&1\\ 1&0&1&0\\ 0&1&0&1\\ 1&0&1&0\\ 0&1&0&0\\ 1&0&0&0\\ 1&1&0&0\end{bmatrix},H_{TS(9,2)}=\begin{bmatrix}1&1&1&1&0&0&0&0&0\\ 0&1&0&1&0&0&0&0&0\\ 1&0&1&0&0&0&0&0&0\\ 1&1&1&1&0&0&0&0&0\\ 0&1&0&1&0&0&0&0&0\\ 1&0&1&0&1&0&0&0&0\\ 1&0&0&0&0&0&1&0&0\\ 0&1&0&0&0&0&0&0&0\\ 0&0&0&0&1&1&0&0&0\\ 1&0&0&0&0&0&1&0&0\\ 0&1&0&0&0&0&0&0&1\\ 0&0&0&0&0&1&0&1&0\\ 1&1&0&0&0&0&0&0&0\\ 0&1&0&0&0&0&0&0&1\\ 1&0&0&0&0&0&0&1&0\end{bmatrix}$. (3)
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$c_{0}$$c_{1}$$c_{2}$$c_{3}$$c_{4}$
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$c_{0}$$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{10}$$c_{11}$$c_{12}$
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}$$x_{7}$$x_{8}$$c_{0}$$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{10}$$c_{11}$$c_{12}$$c_{13}$$c_{14}$
Figure 3: Graphical representation of TS(4, 2) (left), TS(4, 6) (center) and TS(9, 2) (right)
4.1 Spectral manifestation of trapping sets
[…]
4.2 Ihara–Bass zeta function
For a finite graph $G$ let $\mathcal{B}$ denote the non-backtracking operator. The Ihara–Bass zeta function is defined by $\zeta_{G}(u)=\prod_{[C]}\bigl{(}1-u^{\,\ell(C)}\bigr{)}^{-1},$ where the product runs over equivalence classes $[C]$ of primitive cycles and $\ell(C)$ is the length of a cycle,  [19] . Poles of $\zeta_{G}$ occur at reciprocals of eigenvalues of $\mathcal{B}$; therefore each closed cycle contributes a factor $(1-u^{\ell})^{-1}$. When $G$ contains a trapping set, the associated cycles generate poles in $\zeta_{G}(u)$, which manifest as isolated eigenvalues of both $\mathcal{B}$ and the Bethe–Hessian $H_{\beta,J}$. These poles can be interpreted as $\mathbb{Z}/2$-torsion elements: they correspond to non-trivial $1$-cycles ($\beta_{1}>0$) in the underlying feature manifold. Removing trapping sets eliminates the corresponding poles and reduces the number of low-energy modes.
4.3 Bordism triviality and topological obstructions
Let $G=(V,E)$ be a simple graph and let $\phi:V\rightarrow\mathbb{R}^{d}$ embed its vertices as points of a smooth manifold $M_{G}:=\phi(V)\subset\mathbb{R}^{d}$, [20]. We call $G$ bordism-trivial if $M_{G}$ is null–bordant in the appropriate bordism group, i.e. there exists a compact $(d+1)$-manifold $W$ with $\partial W=M_{G}\qquad(\text{or }\partial W=M_{G}\sqcup(-M_{G})\text{ for the unoriented case}),$ and all characteristic numbers of $M_{G}$ vanish: $w_{k}(M_{G})=0,\;k>0,\qquad p_{j}(M_{G})=0,\;j\geq 1.$ Equivalently $M_{G}]=0$ in the oriented bordism group $\Omega^{SO}_{d}$ (and in $\Omega^{O}_{d}$ when orientation is ignored). For a multi-class problem let $\{M_{i}\}_{i=1}^{K}$ be the manifolds obtained from the feature vectors of each class. A bordism obstruction between classes $i$ and $j$ occurs if the disjoint union $M_{ij}:=M_{i}\sqcup(-M_{j})$ is not null-bordant in $\Omega^{SO}_{d}$ (or $\Omega^{O}_{d}$). A convenient detection criterion is the non-vanishing of a characteristic number, e.g. $\exists\,k>0:\;w_{k}[M_{ij}]\neq 0\quad\text{or}\quad\exists\,j:\;p_{j}[M_{ij}]\neq 0.$ If such an obstruction exists for any pair of classes, no graph-based embedding can achieve perfect separation in the high-dimensional feature space, [21].
4.4 Higher signatures and KK–theory
Given a cell complex $\mathcal{M}_{G}$ derived from the graph, let $L_{k}=d_{k-1}^{*}d_{k-1}+d_{k}d_{k}^{*}$ be the combinatorial Hodge Laplacian acting on $k$-cochains, [22]. The kernel dimension satisfies $\dim\ker L_{k}=\beta_{k},$ so that spectral information directly yields Betti numbers.
For any cohomology class $x\in H^{*}(B\pi_{1}(\mathcal{M}_{G});\mathbb{Q})$, the higher signature $\sigma_{x}(\mathcal{M}_{G})=\bigl{\langle}L(\mathcal{M}_{G})\smile x,\,[\mathcal{M}_{G}]\bigr{\rangle}$ pairs the Hirzebruch $L$-class with the fundamental class,  [23]. By the Novikov conjecture, $\sigma_{x}\!\!\pmod{2}$ detects non-trivial $\mathbb{Z}/2$-torsion, hence odd-length trapping cycles render some higher signature non-zero.
The Dirac operator on the bipartite graph, $\mathcal{D}=\begin{pmatrix}0&A_{\!vn}\\ A_{\!vn}^{\!\top}&0\end{pmatrix}$, has spectrum symmetric about zero, [24, 25]. Its analytic index equals the difference of dimensions of positive and negative eigenspaces, which, by the Atiyah–Singer index theorem, coincides with $\beta_{0}-\beta_{1}$, [26]. Thus a non-zero index signals the presence of topological defects (e.g. odd cycles).
In $K$-theory, the Kasparov matrix $D_{\mathrm{Kas}}=\begin{pmatrix}0&S&T\\ S^{\!\top}&0&0\\ T^{\!\top}&0&0\end{pmatrix}$, built from suitable incidence matrices $S$ and $T$, satisfies: $K_{0}=\dim\ker D_{\mathrm{Kas}}^{2}$ (counts zero-modes, i.e. $\beta_{0}$); $K_{1}=\operatorname{rank}D_{\mathrm{Kas}}\bmod 2$ (detects $\mathbb{Z}/2$ torsion, equivalent to the Kervaire invariant), [27] . Non-trivial $K$-groups therefore provide a concise algebraic signature of the topological obstruction introduced by a trapping set, [28].
The continuous genus is a scalar number that captures how “twisted’’ or “curved’’ a graph $G$ (or, more generally, a network built from points and edges) is, by looking at the spectrum of a matrix derived from the graph $\widehat{A}(H)=\frac{1}{2\sqrt{n_{V}}}\Biggl{(}\sum_{\lambda_{i}\in\Lambda^{+}}\!\sqrt{\lambda_{i}}-\sum_{\lambda_{j}\in\Lambda^{-}}\!\sqrt{-\lambda_{j}}\Biggr{)}$, where $n_{V}$ are number of vertices in the graph, [29]. The experimental study of how topological invariants influence spectral embedding is presented in Sec. 5.1.
4.5 Permanent and Bethe–permanent bounds
The permanent of an $m\times m$ matrix $\mathbf{B}=[b_{j,i}]_{j,i}$ over a commutative ring is $\operatorname{perm}(\mathbf{B})=\sum_{\sigma\in S_{m}}\;\prod_{j=1}^{m}b_{\,j,\sigma(j)},$ i.e. the determinant without the sign factor $\operatorname{sgn}(\sigma)$. Exact evaluation of (4.5) is #P-complete, consequently a number of approximation and fast-evaluation schemes have been proposed [30]. The minimum Hamming distance $d_{\min}$ of a QC-LDPC code can be bounded by permanents of its weight matrix [31]. For a parity-check matrix $H(x)$ let $\mathbf{A}=\operatorname{wt}(H(x))$ denote the entrywise Hamming weight. The code distance upper bound, $TS(a=d_{\min},0)$:
$d_{\min}\;\leq\;\operatorname{min}^{*}_{\begin{subarray}{c}\mathcal{S}\subset[h]\\ |\mathcal{S}|=v+1\end{subarray}}\;\sum_{i\in\mathcal{S}}\operatorname{perm}\!\bigl{(}\mathbf{A}_{\mathcal{S}\setminus i}\bigr{)},$ (4)
where $\operatorname{min}^{*}$ denotes the minimum over all $(v+1)$-subsets. Replacing the permanent by its Bethe approximation yields an upper bound on the minimum pseudoweight, which controls the harmfulness of $TS(a,b>0)$,  [32]. The Bethe-permanent of a non-negative matrix $\mathbf{B}$ is defined as $\operatorname{perm}_{\!B}(\mathbf{B})=\exp\!\Bigl{(}-\min_{q\in\mathcal{Q}}\;F_{\rm Bethe}(q)\Bigr{)}$, where $F_{\rm Bethe}$ is the Bethe free energy and $\mathcal{Q}$ denotes the set of factorised marginals, [33]. In our construction we maximize weight for both bounds during protograph and parity-check optimization using EMD/ACE maximization, (Alg. 2, [2]), thereby suppressing low-weight TS and improving the Bethe–Hessian spectrum, Eq. 5.
Figure 4: Dependence of the smallest eigenvalue $\lambda_{\min}$ on the temperature parameter $\beta$. The red curve shows the polynomial part, the blue curve the tail, and the black dashed line is a quadratic approximation to the polynomial part. The coefficient of determination for the fit is $R^{2}=0.9998$.
4.6 Quadratic–Newton estimation of the Nishimori temperature $\beta_{N}$
The root condition (1) can be solved efficiently because the map $\beta\mapsto\lambda_{\min}(\beta)$ is smooth and monotone, Fig. 4. Algorithm 1 describes proposed $\beta_{N}$ estimator. The algorithm typically converges within $\sim 9$ Arnoldi calls, a 6-fold reduction compared with classical bisection [1, 2].
Require : AA (adjacency), $D$ (degree matrix), lower bound $\beta_{\ell}$, upper bound $\beta_{u}$.
1 $i\leftarrow 0$;
2 repeat
3  convergence
4until if $i=0$ then
5 Choose three trial points $\beta_{1}=\beta_{\ell},\;\beta_{2}=(\beta_{\ell}+\beta_{u})/2,\;\beta_{3}=\beta_{u}$;
6
7else
8 Center the interval around the latest estimate $\tilde{\beta}$ and set $(\beta_{1},\beta_{2},\beta_{3})$ accordingly;
9
10for $k=1,2,3$ do
11 Compute $B_{\beta_{k}}=(\beta_{k}^{2}-1)I-\beta_{k}A+D$;
12  Evaluate $\lambda_{k}=\lambda_{\min}(B_{\beta_{k}})$ via Arnoldi;
13
14Fit a quadratic polynomial $p(\beta)=a\beta^{2}+b\beta+c$ to $(\beta_{k},\lambda_{k})_{k=1}^{3}$;
15 Compute the positive root $\tilde{\beta}=\frac{-b-\sqrt{b^{2}-4ac}}{2a}$;
16 if $|\lambda_{\min}(B_{\tilde{\beta}})|<\varepsilon$ then
17 return $\beta_{N}:=\tilde{\beta}$;
18
19else
[…]
21
[…]
Algorithm 1 Quadratic–Newton estimation of the Nishimori temperature $\beta_{N}$
4.7 Graph construction from CNN features
Given an image $x$, the backbone network $f_{\theta}$ (MobileNetV2, [34]) yields a feature vector $\mathbf{h}\in\mathbb{R}^{1280}$. For each class $c$ we retain a subset $\mathcal{I}^{(c)}$ of $s\ll 1280$ most discriminative indices, e.g. those with largest absolute difference between class-conditional means. The reduced feature vector for class $c$ is $\mathbf{z}^{(c)}=(h_{i})_{i\in\mathcal{I}^{(c)}}\in\mathbb{R}^{s}$. A similarity kernel (Cosine, gaussian and etc) builds a weighted adjacency matrix between the $K$ class representatives, (Eq. 8 in [2]):
$A_{cd}=\begin{cases}\exp\!\bigl{(}-\,\gamma\,d_{\mathrm{cos}}(\mathbf{z}^{(c)},\mathbf{z}^{(d)})^{2}\bigr{)},&c\neq d,\\ 0,&c=d,\end{cases}$
where $\gamma$ - kernel bandwidth, $d_{\mathrm{cos}}$ - cosine-distance kernel. To enforce sparsity we retain only the $p$ strongest edges per vertex. The resulting sparse adjacency is then mapped onto a MET QC-LDPC parity-check matrix by selecting an appropriate protograph and CPMs (Section 3).
5 Experimental evaluation
All experiments use the same MobileNetV2 pre-trained backbone, Fig. 5. The feature extractor (CNN layers) is frozen, continuous features are binarized using sign, only the graph embedding (replace of feed-forward MLP) and final classifier are trained. We use following datasets: ImageNet10: 10-class subset (13 000 images, 10 00010\,000 training, 3 0003\,000 testing). ImageNet100: 100-class subset (130 000130\,000 training, 5 0005\,000 testing; 1 3001\,300 training and $50$ test samples per class). The ImageNet10 and ImageNet10 use 32-D spectral and 64-D respectively.
Figure 5: Pipeline: MobileNetV2 feature extraction followed by graph spectral embedding.
5.1 Influence of Topological Invariants on Spectral Embedding
For the purpose of illustration we consider six frequently occurring trapping sets $TS(4,2)$, $TS(4,6)$, $TS(9,2)$, $TS(13,6)$, $TS(26,20)$, $TS(28,22)$. Table 1 collects the most relevant topological and spectral quantities for each of them. From a spectral-embedding perspective a trapping set can be viewed as a locally dense cluster of edges that is attached to only a few low-degree check nodes. Such a configuration produces negative eigenvalues of the Bethe–Hessian matrix $H_{r}$ (see the column “# neg. $\lambda(H_{1})$”) and, more importantly, gives rise to low-energy modes that are poorly separated by any embedding based on the graph Laplacian. Consequently, graphs containing many of these substructures generate highly correlated features, which in turn can dramatically deteriorate the performance of downstream classifiers.
Table 1: Topological and spectral invariants of Trapping sets.
Invariant $TS(4,2)$ $TS(4,6)$ $TS(26,20)$ $TS(9,2)$ $TS(13,6)$ $TS(28,22)$
$\rho$ (spectral radius) 1.618 4.000 2.7545 8.9168 10.4154 13.5644
$r_{\rm crit}=\sqrt{\rho}$ 1.272 2.000 1.6597 2.9861 3.2273 3.6830
$\{\lambda(H_{1})<0\}$ (negative modes at $r=1$) 1 2 1 1 0 0
$\widehat{A}$ (continuous genus) 1.007 1.529 3.5896 3.0687 4.0313 7.2670
$K_{0}$ (zero-dimensional Betti number) 1 5 7 12 17 45
$K_{1}$ (one-dimensional Betti number) 1 1 1 0 1 1
Kervaire invariant $\kappa$ 1 1 1 0 0 0
$w_{2}$ (Stiefel–Whitney) 1 1 1 1 1 1
Mod-2 Betti $(\beta_{0},\beta_{1})$ (1,0) (1,0) (1,0) (1,0) (1,0) (1,0)
Bordism obstruction 0 0 0 0 0 1
To assess the practical impact of these invariants we built two graphs that serve as adjacency matrices in a convolutional architecture trained on the ImageNet-100 benchmark. Optimized irregular MET QC-LDPC toric graph. The protograph matrix (3.6 average column weight) is
$M(H)=\begin{bmatrix}2&2&0&0&0\\ 2&2&1&0&0\\ 1&0&1&1&0\\ 1&0&0&1&1\\ 2&0&0&0&1\end{bmatrix}$,
lifted with a circulant size of $520$, [2]. During the construction we deliberately eliminated harmful trapping sets; the remaining trapping sets is the relatively benign $TS(26,20)$, which possesses a single negative Bethe–Hessian eigenvalue and a modest continuous genus (see Table 1). After a Bayesian classifier we obtained Top-1 accuracy: $0.6724\pm 2.46\times 10^{-6}$; Top-3 accuracy: $0.8582\pm 8.53\times 10^{-7}$. Thus, with an average degree of only $3.6$ we achieve performance comparable to much denser near regular graphs. Non-optimized irregular MET QC-LDPC toric graph (same $\bar{w}=3.6$) was built without filtering out low-genus, high-mode traps. Consequently it contains the sets $TS(4,2)$, $TS(9,2)$, $TS(13,6)$ and $TS(28,22)$. These introduce several additional negative Bethe–Hessian eigenvalues (up to three for $TS(28,22)$) and raise $\widehat{A}$, thereby generating many low-energy directions that corrupt the spectral embedding. The resulting classification scores are Top-1 accuracy: $0.1368\pm 3.72\times 10^{-4}$; Top-3 accuracy: $0.2450\pm 5.63\times 10^{-5}$. The more than four-fold drop in top-1 performance underscores how a few detrimental trapping sets can cripple the discriminative power of graph-based features. For reference we generated dense Erdős–Rényi graphs (no QC structure, original weight 5 and 10) with average degrees $d\approx 3$ and $d\approx 5$ (almost regular) after cycle broken. After six independent runs the observed accuracies were top-1 $0.617467\pm 3.72E-05$ and top-3 $0.8283\pm 2.54E-05$ (for weight 5) and top-1 $0.659333\pm 1.30E-05$ and top-3 $0.85396667\pm 8.4E-06$ (for weight 10). Even though these dense random graphs outperform the non-optimised QC-LDPC design, they are considerably less efficient (higher degree, no structured protection) and lack the flexibility offered by quasi-cyclic constructions. Negative Bethe–Hessian modes identify substructures that generate low-energy directions in the spectral embedding; each such direction reduces feature orthogonality. The continuous genus $\widehat{A}$ provides a quantitative measure of how “dangerous’’ a trap is: larger values correlate with more severe degradation of downstream performance. By pruning high-genus, multi-mode traps (e.g. $TS(4,2)$, $TS(28,22)$) we can keep the graph sparse ($\bar{w}\approx 3.6$) while retaining classification accuracy comparable to much denser Erdős–Rényi graphs, and improve top-3 accuracy important for graph ensemble (using unequal protection of features, introduce high degree columns). The presented design methodology—selecting protographs, eliminating harmful trapping sets, and fine-tuning via Bayesian optimization—yields QC-LDPC graphs that combine structured protection, and strong discriminative power. All auxiliary material (list of small weight TS, TS submatrix, parity-check matrix, eigenvalue trajectories, topology invariants and related quadratic form signatures, source code, etc) is available in the public repository [35].
Figure 6: Spectral embedding histogram with adjacency matrix of clusters (visualized by colors) under estimated Nishimori temperature $\beta_{N}$. (left) MET QC-LDPC Torical graph. (right) QC Spherical graph.
5.2 Graph Ensembles for ImageNet-100
We construct two families of MET–QC-LDPC graphs and then combine them into an ensemble that produces a final class decision by majority voting, Fig. 6. Spherical graphs – each instance consists of a single circulant ring, has a fixed column weight $k=10$ and CPM size $L=2600$. Toroidal graphs – each instance is built from two independent circulant rings (forming a torus), uses $L=520$, and exhibits an average column weight of 3.6. Both families are first optimised with respect to the permanent bound (4) and the Bethe-permanent bound [27]. This optimisation explicitly suppresses low-weight trapping sets that would otherwise degrade the spectral embedding. From the two families we generate three affinity graphs (see Fig. 6), feed each graph into an identical classifier and obtain three independent predictions: $\widehat{y}_{1},\;\widehat{y}_{2},\;\widehat{y}_{3}.$ The final label is obtained by a simple majority vote. In the rare case that the three votes do not agree we invoke an arbiter network to resolve the conflict. Formally,
$\widehat{y}_{\text{res}}=\begin{cases}\widehat{y}_{1},&\text{if }\widehat{y}_{1}=\widehat{y}_{2}=\widehat{y}_{3},\\ \operatorname{Arb}\!\bigl{(}\widehat{y}_{1},\widehat{y}_{2},\widehat{y}_{3}\bigr{)},&\text{otherwise},\end{cases}$
where $\operatorname{Arb}(\cdot)$ denotes the output of the arbiter network. Instead of hard voting we can average the class-wise posterior probabilities supplied by each graph $p_{c}=\frac{1}{3}\sum_{i=1}^{3}p^{(i)}_{c},\widehat{y}_{\text{soft}}=\arg\max_{c}p_{c},$ with $p^{(i)}_{c}$ being the probability of class $c$ estimated by the $i$-th graph. The soft-voting rule often yields smoother confidence scores while preserving the benefits of ensemble diversity.
Table 2: Classification results on ImageNet10 using graph embeddings.
Class Precision Recall F1-Score Support
0 1.00 0.98 0.99 300
1 0.99 0.99 0.99 300
2 1.00 0.99 0.99 300
3 0.99 0.94 0.96 300
4 0.97 0.97 0.97 300
5 0.99 0.99 0.99 300
6 0.94 0.95 0.95 300
7 0.98 0.98 0.98 300
8 0.94 0.98 0.96 300
9 0.96 1.00 0.98 300
Aggregate Metrics
Accuracy 0.98 3000
Macro Avg 0.98 0.98 0.98 3000
Weighted Avg 0.98 0.98 0.98 3000
5.3 Training protocol
For each class we compute $\mathbf{z}^{(c)}$, build a sparse similarity graph, embed it via the Bethe–Hessian at the estimated Nishimori temperature $\beta_{N}$ (Algorithm 1), and extract the $r$ eigenvectors ($r=32$ for ImageNet10, $r=64$ for ImageNet100), Fig. 6 . These eigenvectors constitute a low-dimensional embedding $\mathbf{e}_{i}$ for every training sample. A linear classifier (softmax) is trained on $\{\mathbf{e}_{i}\}$ using cross-entropy loss.
Figure 7: (Left) Confusion matrix heatmap for ImageNet100. (Right) Per-class top-1 accuracy heatmap.
Figure 8: Spectral embedding histogram for ’cock’ and ’hen’ classes
5.4 Results
Table 2 reports top-1 accuracy, precision, recall and $F_{1}$ scores for ImageNet10. Spherical Graph Spectral Embedding achieves $98.7\%$ top-1 accuracy on ImageNet10, while Fig. 7 shows the confusion matrix for ImageNet100. The mixed ensemble combining spherical and toroidal graphs through majority voting yields the best performance. For ten classes, discriminative features rarely overlap, but scaling to one hundred classes causes many informative dimensions to intersect, reducing accuracy for both plain CNNs and spectral classifiers. Our experiments identified optimal graphs for ImageNet100 consisting of a spherical adjacency with weight $k=10$ and circulant size $L=2600$, paired with a toroidal graph having weight $k=10$, protograph size $26\times 26$ (containing 5 non-zero MET CPMs of weight 2), and CPM size $L=100$, as visualized in Fig. 6. This configuration delivers $77.84\%$ top-1 accuracy when processed through the spectral pipeline. To address frequent confusion between the two highest class probabilities, we introduced a lightweight MLP arbitrator trained on the same embeddings. The arbitrator achieves $98.43\%$ average pairwise accuracy on training data, though the most challenging pairs remain at approximately $67\%$ accuracy. During inference, the system examines the three largest spectral scores $\{p_{(1)},p_{(2)},p_{(3)}\}$, invoking the arbitrator when the margin $p_{(1)}-p_{(2)}$ falls below a preset threshold $T$. This uncertainty-driven fallback mechanism boosts top-1 accuracy from $77.84\%$ to $\mathbf{82.73\%}$ while maintaining the original runtime efficiency and memory footprint of the single-graph approach. The Erdős–Rényi graph spectral embedding demonstrates $69.1\%\pm 2.88\times 10^{-5}$ precision under Nishimori temperature across five trials, averaged over all random seeds. Our three-graph ensemble achieves $\mathbf{82.7\%}$ accuracy on ImageNet100 while using just one twentieth of the original feature dimension and memory footprint, representing a $13.6\%$ improvement over standard Erdős–Rényi graphs without large weights. Classification challenges emerge most prominently between class 7 ("cock") and class 8 ("hen"), as shown in Fig. 7, with training accuracy reaching $85\%$ but validation accuracy dropping to $63\%$. The spectral embedding histogram under Nishimori temperature (Fig. 8) reveals Gaussian intersections with heavy tails in this problematic region. Another challenging area occurs around class 44 ("spider"), where we observe three-class overlap in the heatmap visualization, Fig. 7.
6 Future Directions
The present pipeline keeps the CNN backbone frozen, employs a similarity kernel and uses a MET-QC-LDPC adjacency. A natural next step is to make all three components differentiable and train them jointly. One can back-propagate through the feature extractor $f_{\theta}$, learn a Mahalanobis (or other) metric for the similarity kernel, and adjust the CPM that define the QC-LDPC parity-check matrix by gradient descent or reinforcement learning. Rich multimodal embeddings such as contrastive language-image pre-training provide powerful initializations for this end-to-end scheme. Embedding topological information directly into the loss is another promising avenue. TS correspond to unwanted $\mathbb{Z}/2$-torsion and low-genus cycles, which appear as negative eigenvalues of the Bethe–Hessian. Adding a differentiable penalty proportional to the number (or an estimator) of such negative modes—e.g. could suppress TS during training instead of removing them post-hoc. Extending experiments from ImageNet-100 to the ImageNet-10k benchmark, applying the method to video streams could reveal new ways of handling high-dimensional structured data while preserving a tiny memory footprint suitable for edge devices. The sparse adjacency matrices produced by our design are attractive priors for GNN. Incorporating them into message-passing GCN/GAT layers, using them to sparsify self-attention in Transformers, or employing diffusion-based graph convolutions may combine the interpretability of physics-inspired graphs with the expressive power of learned GNNs.
7 Conclusions
We introduced a physics-inspired embedding that maps CNN features onto spins of a Random-Bond Ising Model defined on MET-QC-LDPC graphs. Trapping sets create Ihara–Bass zeta poles that appear as isolated eigenvalues of the non-backtracking operator, encoding $\mathbb{Z}/2$-torsion, Betti numbers and bordism obstructions. A Nishimori-temperature estimator evaluates the Bethe–Hessian eigenvalue at three temperatures, fits a quadratic interpolant and applies one Newton step, reaching convergence in $\approx 9$ Arnoldi iterations—six times faster than bisection. Graph design combines permanent and Bethe-permanent bounds with ACE/EMD to suppress trapping sets, yielding sparse spherical/toridal graphs. These compress MobileNetV2’s 1280-dimensional features to 32 dimensions for ImageNet-10 and 64 for ImageNet-100, achieving top-1 accuracies of 98.7% and 82.7% while cutting parameters and memory by factors of 40 and 20. The proposed method unites statistical mechanics, coding theory and topology into a sparse QC-LDPC embedding that exceeds baselines with lower cost, enabling efficient DNN.
Appendix A Trapping sets as topological defects in the spectral embedding
In this appendix, we give a completely rigorous step-by-step proof that any trapping-set subgraph $\mathcal{T}$ of a QCLDPC Tanner graph creates a low-energy (negative) mode of the Bethe–Hessian $H_{\beta,J}$ and, equivalently, produces a pole of the Ihara–Bass zeta function. The argument is split into four formal statements.
Notation
  • $\mathcal{G}=(V,E)$ – undirected simple graph, $|V|=n$, $|E|=m$.
  • $H\in\{0,1\}^{m\times n}$ – binary incidence (parity-check) matrix of a QC-LDPC code; rows = check nodes, columns = variable nodes.
  • For each edge $\{i,j\}\in E$ we write $J_{ij}\in\mathbb{R}$ for the Ising coupling.
  • For $\beta>0$ the Bethe–Hessian is
    $H_{\beta,J}=\operatorname{diag}\!\Bigl{(}1+\sum_{k\in\partial i}\frac{\tanh^{2}(\beta J_{ik})}{1-\tanh^{2}(\beta J_{ik})}\Bigr{)}-\Bigl{[}\frac{\tanh(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}\Bigr{]}_{i\neq j}.$ (5)
  • $Q_{\beta,J}(x)=x^{\!\top}H_{\beta,J}x$ is the associated quadratic form.
  • A trapping set $\mathcal{T}=(V_{\mathcal{T}},E_{\mathcal{T}})$ is the subgraph induced by a subset $V_{\mathcal{T}}\subseteq V$ together with all incident check nodes. Its incidence matrix is denoted $H_{\mathcal{T}}$.
  • $H_{\beta,J}^{\mathcal{T}}$ denotes the principal submatrix of $H_{\beta,J}$ obtained by deleting rows and columns indexed by $V\setminus V_{\mathcal{T}}$.
Step 1 – The restricted Bethe–Hessian is a signed Laplacian
Lemma A.1.
Let $\mathcal{T}$ be any trapping set. Then $H_{\beta,J}^{\mathcal{T}}=D^{\mathcal{T}}(\beta)-B^{\mathcal{T}}(\beta),$ where
$B_{ij}^{\mathcal{T}}(\beta)$ $=\frac{\tanh(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})},$ $i\neq j,\;(i,j)\in E_{\mathcal{T}},\$4pt]D^{\mathcal{T}}_{ii}(\beta)$ $=1+\sum_{j:(i,j)\in E_{\mathcal{T}}}\frac{\tanh^{2}(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}.$
Thus $H_{\beta,J}^{\mathcal{T}}$ is the (weighted) signed Laplacian of the subgraph $\mathcal{T}$.
Proof.
Inspect (5). For any vertex $i\in V_{\mathcal{T}}$ the diagonal entry equals precisely $D^{\mathcal{T}}_{ii}(\beta)$. If $(i,j)\in E_{\mathcal{T}}$ and $i\neq j$, the off–diagonal entry is $\frac{\tanh(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}=-B_{ij}^{\mathcal{T}}(\beta)$. All entries that correspond to vertices outside $V_{\mathcal{T}}$ are removed when we take the principal submatrix, which yields the claimed decomposition. ∎
Step 2 – One negative eigenvalue per simple cycle at Nishimori temperature
We first recall the Nishimori condition.
Definition 1 (Nishimori temperature, [1, 2, 3]).
Let $\beta_{N}>0$ be such that for every edge belonging to a simple cycle of $\mathcal{T}$
$\tanh^{2}(\beta_{N}J_{ij})=1.$ (6)
The inverse temperature $\beta=\beta_{N}$ is called the Nishimori temperature.
Proposition 1.
Let $C\subseteq\mathcal{T}$ be a simple cycle of length $\ell\geq 2$ satisfying (6) on each of its edges. Then the restriction $H_{\beta_{N},J}^{C}$ possesses exactly one negative eigenvalue; all other eigenvalues are non-negative. The corresponding (up to a gauge transformation) eigenvector is $v=(1,-1,1,-1,\dots)\in\mathbb{R}^{\ell},$ i.e. it alternates sign along the cycle.
Proof.
Because the Tanner graph of an LDPC code is bipartite, every cycle $C$ is even; write $\ell=2r$. Hence a gauge transformation $x_{i}\;\mapsto\;s_{i}\,x_{i},\qquad s_{i}\in\{\pm 1\},$ can be chosen so that after the change of variables every edge of $C$ has positive coupling: $\tanh(\beta J_{ij})=+\sqrt{1-\varepsilon},\qquad\varepsilon:=1-\tanh^{2}(\beta J_{ij})>0.$ The parameter $\varepsilon$ depends on the edge, but for all edges of $C$ we have $\varepsilon\downarrow 0$ as $\beta\to\beta_{N}$. For an edge $(i,j)\in C$ we compute $\frac{\tanh^{2}(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}=\frac{1-\varepsilon}{\varepsilon},\qquad\frac{\tanh(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}=\frac{\sqrt{1-\varepsilon}}{\varepsilon}.$ All other edges (those not belonging to $C$) contribute bounded terms that remain finite as $\varepsilon\to 0$; we denote their total contribution by a matrix $R_{\varepsilon}$ satisfying $R_{\varepsilon}=O(1)$. Let $A_{C}$ be the adjacency matrix of the cycle (entries equal to 1 on the two neighboring positions, zero otherwise). Using Lemma A.1 we obtain for the restriction to $C$: $H_{\beta,J}^{C}=\Bigl{(}1+\frac{2(1-\varepsilon)}{\varepsilon}\Bigr{)}I-\frac{\sqrt{1-\varepsilon}}{\varepsilon}A_{C}+R_{\varepsilon}.$ Factorizing $1/\varepsilon$ and expanding $\sqrt{1-\varepsilon}=1+O(\varepsilon)$ gives the asymptotic expansion
$H_{\beta,J}^{C}=\frac{1}{\varepsilon}\bigl{(}2I-A_{C}\bigr{)}-I+O(\varepsilon).$ (7)
The matrix $L_{C}:=2I-A_{C}$ is precisely the (combinatorial) Laplacian of a cycle. Its eigenvalues are well known: $\mu_{k}=2-2\cos\!\Bigl{(}\frac{2\pi k}{\ell}\Bigr{)},\qquad k=0,\dots,\ell-1.$ The associated orthonormal eigenvectors are the discrete Fourier modes. In particular, $\mu_{0}=0,\qquad\mu_{r}=4\quad(\text{since }\ell=2r),$ and $\mu_{k}>0$ for all $k\neq 0$. Applying 7 to an eigenpair $(\mu_{k},u^{(k)})$ of $L_{C}$ yields $H_{\beta,J}^{C}u^{(k)}=\Bigl{(}\frac{\mu_{k}}{\varepsilon}-1+O(\varepsilon)\Bigr{)}u^{(k)}.$ Hence the eigenvalues of $H_{\beta,J}^{C}$ behave as
$\lambda_{k}(\varepsilon)=\frac{\mu_{k}}{\varepsilon}-1+O(\varepsilon).$ (8)
Because $\mu_{0}=0$, 8 gives $\lambda_{0}(\varepsilon)\to-1$, whereas for every $k\neq 0$ we have $\mu_{k}>0$ and thus $\lambda_{k}(\varepsilon)\to+\infty$ as $\varepsilon\downarrow 0$. Consequently, in the limit $\beta\to\beta_{N}$ there is exactly one negative eigenvalue (tending to 1-1) and all remaining eigenvalues are non-negative. The eigenvector belonging to $\mu_{0}=0$ is the constant vector in the gauge- transformed basis. Undoing the gauge transformation multiplies each component by the sign factor $s_{i}$, which yields precisely the alternating pattern $v=(1,-1,1,-1,\dots)$. Since the cycle length is even, this vector is well defined. ∎
Step 3 – The number of negative eigenvalues equals the first Betti number
Recall that for a finite graph $\Gamma$ the first Betti number (the dimension of its cycle space) is $\beta_{1}(\Gamma)=|E(\Gamma)|-|V(\Gamma)|+c(\Gamma),$ where $c(\Gamma)$ denotes the number of connected components.
Theorem A.2.
Let $\mathcal{T}$ be a trapping set and assume that every edge belonging to a simple cycle of $\mathcal{T}$ satisfies the Nishimori condition (6). Then the number $q_{\mathcal{T}}$ of negative eigenvalues of the restricted Bethe–Hessian equals its first Betti number: $q_{\mathcal{T}}=\beta_{1}(\mathcal{T}).$
Proof.
Decompose $\mathcal{T}$ into its $c(\mathcal{T})$ connected components $\mathcal{T}^{(1)},\dots,\mathcal{T}^{(c)}$. Because $H_{\beta_{N},J}^{\mathcal{T}}$ is block-diagonal with respect to this decomposition, the total number of negative eigenvalues is the sum over the components. Fix a component $\mathcal{C}$. Choose a spanning forest $F$ of $\mathcal{C}$; the edges of $F$ form a tree and thus contribute no negative eigenvalue (its Bethe–Hessian reduces to a positive semidefinite matrix, see Lemma A.1). Every edge $e\in E(\mathcal{C})\setminus F$ creates exactly one independent cycle when added to $F$. By Proposition 1 each such cycle produces one negative eigenvalue of $H_{\beta_{N},J}^{\mathcal{C}}$, and different edges generate linearly independent cycles, hence distinct negative eigenvalues. Therefore, $q_{\mathcal{C}}=|E(\mathcal{C})|-|V(\mathcal{C})|+1.$ Summing over all components gives $q_{\mathcal{T}}=\sum_{c=1}^{c(\mathcal{T})}\!\bigl{(}|E(\mathcal{T}^{(c)})|-|V(\mathcal{T}^{(c)})|+1\bigr{)}=|E(\mathcal{T})|-|V(\mathcal{T})|+c(\mathcal{T})=\beta_{1}(\mathcal{T}),$ which is precisely the claimed equality. ∎
Step 4 – Negative eigenvalues give poles of the Ihara–Bass zeta function
We recall the non-backtracking (Hashimoto) matrix and the Ihara–Bass zeta function.
Definition 2 (Ihara–Bass zeta function).
For a finite graph $\Gamma$ let $\mathcal{B}(\Gamma)$ be its non-backtracking operator. The Ihara–Bass zeta function is $\zeta_{\Gamma}(u)=\prod_{[C]}\bigl{(}1-u^{\,\ell(C)}\bigr{)}^{-1},$ where the product runs over all primitive (non-repeating) cycles $[C]$ and $\ell(C)$ denotes their length.
A classical identity (see e.g. [7]) links $\zeta_{\Gamma}$, the non-backtracking matrix, and the Bethe–Hessian:
$\zeta_{\Gamma}(u)=\frac{1}{\det(I-u\,\mathcal{B})}=\bigl{(}1-u^{2}\bigr{)}^{-|E|+|V|}\det\!\bigl{(}H_{\beta,J(u)}\bigr{)}^{-1},$ (9)
with the bijection $u=\tanh(\beta J)$ between the spectral parameter $u$ and the temperature $\beta$.
Corollary A.2.1.
Let $\Gamma$ be any graph. If for some $u_{c}$ we have $\det(H_{\beta,J(u_{c})})=0$, then $u_{c}$ is a pole of the Ihara–Bass zeta function. In particular, every negative eigenvalue identified in Theorem A.2 produces a pole at $u_{c}=\tanh(\beta_{N}J_{e})$, where $J_{e}$ is the coupling on the corresponding cycle edge.
Proof.
Equation (9) shows that $\zeta_{\Gamma}(u)$ diverges exactly when $\det(H_{\beta,J(u)})=0$. By Theorem A.2, at the Nishimori temperature each independent cycle forces an eigenvalue of $H_{\beta_{N},J}$ to cross zero (from negative for $\beta>\beta_{N}$ to positive for $\beta<\beta_{N}$). Hence the corresponding $u_{c}=\tanh(\beta_{N}J)$ is a pole of $\zeta_{\Gamma}(u)$. ∎
The exact correspondence proved above relies on the strict Nishimori condition $\tanh^{2}(\beta_{N}J_{ij})=1$ for every edge belonging to a cycle. In practice this holds only for uniform couplings $|J_{ij}|=\infty$ or when the distribution of $J$ is chosen so that the Nishimori line coincides with $\tanh^{2}=1$. Nevertheless, the mechanism is robust: Even if $\tanh^{2}(\beta J_{ij})<1$ but close to 1 on a cycle, the two terms in (5) become large with opposite sign. The matrix $H_{\beta,J}^{C}$ then has one very small eigenvalue, which becomes negative as soon as $\beta$ exceeds a critical value that is close to $\beta_{N}$. Consequently, away from the exact Nishimori line the number of negative eigenvalues need not equal $\beta_{1}(\mathcal{T})$, but it is still bounded below by the number of balanced even cycles and above by the total number of independent cycles. In other words, $\beta_{1}(\mathcal{T})-\delta\leq q_{\mathcal{T}}\leq\beta_{1}(\mathcal{T})$ for a small defect $\delta$ that vanishes as the Nishimori condition is approached. The Ihara–Bass identity (9) holds for any graph, irrespective of the values of $J$. Therefore trapping sets always generate poles of $\zeta_{\Gamma}$; at the exact Nishimori temperature these poles correspond to eigenvalue crossings of the Bethe–Hessian. Thus the result proved in Steps 1–4 is a sharp statement for the idealized Nishimori point, while the qualitative picture (cycles $\rightarrow$ low-energy modes $\rightarrow$ zeta-function poles) persists under realistic, approximate conditions.
Conclusion
By Lemma A.1, Proposition 1, Theorem A.2 and Corollary A.2.1 we obtain the exact correspondence
[…]
Hence the signature of the Bethe–Hessian
$\operatorname{sig}\!\bigl{(}H_{\beta_{N},J}\bigr{)}=(p,q,r)$
[…]
Let $q$ denote the number of negative eigenvalues of $H_{\beta_{N},J}$. Theorem A.2 shows that each such eigenvalue is in one-to-one correspondence with an independent cycle (i.e. a trapping set) in the graph. Deleting every induced subgraph whose edge set contributes to $q$ forces the Bethe–Hessian to become positive semidefinite, $H_{\beta_{N},J}\;\succeq\;0.$ Because of the Ihara–Bass identity (9), positivity of $H_{\beta_{N},J}$ implies that its determinant never vanishes; consequently all poles of the zeta function disappear. Moreover, the eigenvectors associated with the negative eigenvalues constitute low-energy directions in the spectral embedding; their removal yields a clean, well-conditioned embedding whose geometry is no longer corrupted by spurious modes, thereby improving class separability. In coding-theoretic language these low-energy modes are precisely the topological defects (even cycles with odd degree nodes) that generate pseudocodewords: spin configurations that lie in wide basins of attraction around genuine codewords. When such pseudo-codewords are present, a Bethe–Hessian based training algorithm can become trapped in these basins, leading to sub-optimal solutions and reduced discriminative power for downstream classifiers. By excising the defect-producing subgraphs we simultaneously achieve
$H_{\beta_{N},J}\;\succeq 0,\qquad\zeta_{\mathcal{G}}(u)\text{ pole\textendash free},$
and we remove the pseudocodeword basins, thereby restoring discriminative capacity to the spectral embedding.
From a different perspective, a TS shows up as a negative mode of $H_{\beta,J}$, this mode creates a wide basin of attraction in the underlying energy landscape. Wide basins are precisely what allows an LDPC ensemble to approach Shannon capacity, because they keep the iterative decoder inside a region of locally stable configurations long enough for message passing to converge. The weight-distribution analysis [36] (and its natural extension to replica-symmetry-breaking [37]) shows that, when the degree-distribution pair satisfies the flatness (stability) condition required for capacity on the binary erasure channel – and, by continuity, also for the AWGN case treated with the Bethe–Hessian – the minimum distance can grow only logarithmically with the block length. Hence an ensemble that contains codewords of very large weight (i.e. a trapping set $\mathrm{TS}(a,0)$ with $a\gg\log n$) cannot be capacity-achieving.
Conversely, ensembles enriched with many pseudo-codeword trapping sets, $\mathrm{TS}(a,b\neq 0)$, possess the low-energy negative modes required for those wide basins and can therefore operate arbitrarily close to capacity. The following balance is therefore essential: Modest-weight codewords ($b=0$) guarantee a non-trivial minimum distance, preventing the code from collapsing into a completely dense set of low-weight words; Non-zero-bias trapping sets ($b\neq 0$) generate the wide basins that make iterative decoding robust and enable capacity achievement.
This trade-off creates a controlled imperfection in the graph: it limits the asymptotic growth of $d_{\min}$ while simultaneously improving the discriminative power of the spectral embedding used for classification.
Exactly the same compromise appears in modern learning systems. In paper  [38] introduced the notion of a robust ensemble (RE) – a dense region of weight configurations that can be accessed by interacting replicas (e.g. focusing Belief Propagation, fBP). Their analysis shows that algorithms which bias the dynamics toward such dense regions avoid narrow, poor-performing minima and instead settle in wide, high-entropy basins. These REs are the algorithmic counterpart of the negative Bethe–Hessian modes we identify as trapping sets.
[…]
CONFLICT OF INTEREST
The authors of this work declare that they have no conflicts of interest.